# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

### v1

@@ 4 How are moderators of different communities perceived differently by their communities?
Method. To quantify a community’s health, we divide communities into quartiles based on the responses of their community members to a recent survey assessing the current state of 2,151 communities (Weld, Zhang, and Althoff 2022). Survey responses were collected between May-July 2021, a period overlapping the the end of our mod discourse data time range. For this analysis only, we exclude communities which were not surveyed. Community topic is classified into one of six topical categories using our topic classifier (§3.2). We quantify community size as the volume of submitted content: the average number of posts and comments per day over our study period, which we use to group communities with similar size. For each group, we compute the average fraction of mod discourse, both positive and negative, as well as the total amount of mod discourse.
@@
Results. Different aspects of community health are associated with better and worse perceptions of moderators. The smallest-feeling 25% of communities, based on member self-reports, use 27% more positive and 16% less negative sentiment to discuss their moderators than the largest-feeling 25% of communities (Figure 2a). Figure 2a-e shows that communities which rate themselves as having higher quality content and being more trustworthy, more inclusive, and more safe all use more positive and less negative to discuss their moderators, as well. The difference between the top- and bottom- quartiles are most pronounced for the communities that consider themselves the most- vs. least- trustworthy, with the most-trustworthy communities using 34% more positive and 22% less negative sentiment to describe their moderators (Figure 2b).
Figure 3: Moderators in small communities are perceived more positively and less negatively than moderators in large communities. (a) Tiny communities’ discussion of their moderators has 6.0×6.0$\times$6.0 × more positive sentiment than huge communities, while huge communities further use 2.1×2.1$\times$2.1 × more negative sentiment to describe their mods, relative to the total amount of mod discourse. (b) Tiny communities also tend devote a relatively large proportion of their content to discussing their moderators, with tiny communities having on average 6.1×6.1$\times$6.1 × more posts and comments discussing moderators than larger communities.
@@
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities seem to have more favorable perceptions of their moderators (Figure 3a), with tiny communities with $<10$ posts+comments/day using 6.0×6.0$\times$6.0 × more positive and 0.46×0.46$\times$0.46 × as much negative sentiment to describe their mods than huge communities (those with more than 10k posts and comments per day), relative to the total amount of mod discussion. Tiny communities also devote 7.6×7.6$\times$7.6 × as many total posts and comments to discussing their mods than huge communities, although this effect diminishes with increasing community size—the average amounts of mod discussion in medium, large, and huge communities are all within ±$\pm 7.0\%$ of one another (Figure 3b).
@@
Examining communities of different topics, hobby communities have generally the most positive perception of their moderators, while meme and news-sharing communities have the most negative perception, with meme and news-sharing communities having 46% more of their mod discourse have negative sentiment (Figure 4). Meme and news sharing communities also have more mod discourse than hobby, identity, and photo/video communities. Discussion communities have the most mod discourse of any community topic, on average (Figure 4).
Figure 4: Moderator perceptions varies significantly across communities with different topics. (a) On average, hobby communities have the best perception of their moderators, while meme and news communities have the poorest. (b) Discussion communities have largest amount of mod discourse of any community topic. In this and all other figures, points represent mean estimates alongside bootstrapped 95% confidence intervals.
@@
Implications. Surveys of redditors have shown that they consider a wide range of factors to be important to the overall ‘health’ of their communities (Weld, Zhang, and Althoff 2024). Here we show that communities that are doing well with regards to factors widely considered to be important, such as safety and quality of content, tend to perceive their moderators more positively. Many of these factors are only indirectly controlled by the moderators, for example quality of content—while moderators can set and enforce rules aimed at improve the quality of content in their communities, quality of content is ultimately a function of the content submitted by community members, not moderators. This can lead to moderators being ‘blamed’ for problems largely outside of their control, which has been previously shown in small-n surveys (Matias 2019a; Jhaver, Frey, and Zhang 2023) and is further supported empirically here, as shown in Figure 2.
@@
Small communities appear to both discuss their moderators more, as well as use more positive sentiment in their discussions (Figure 3). Several factors may contribute to this, including that the increased anonymity that comes with participating in a larger community may make people feel more comfortable speaking negatively about the moderators, and that smaller communities are more likely to be newly formed and thus still establishing moderation norms, leading to more mod discourse (Hwang and Foote 2021; Seering and Kairam 2022).
Moderating an online community is a challenging task, with many factors outside the moderators’ control. Some community topics are simply more contentious than others, and therefore more challenging to moderate. We suspect this is a contributing factor to apparent differences in perception of moderators between communities with different topics (Figure 4). To control for this, in subsequent analyses, we validate our findings by stratifying the analyses by community topic and ensuring that trends in results are qualitatively similar, except when stated otherwise.
5 What can moderators do to improve how they are perceived by their community?
@@
Moderators have a great deal of autonomy to run their communities as they see fit, including enforcing rules by removing content and growing the mod team by recruiting or appointing new moderators. Moderators also may (or may not) participate in the community as ‘regular community members’ in addition to their official mod duties. In this section, we examine the impact that different strategies in each of these topics may have on communities’ perceptions of their moderators, and identify promising suggestions for moderators.
Figure 5: Moderators in communities with lower workloads are perceived more positively and less negatively than moderators in communities with high workloads. Communities with lower moderator workloads (more moderators relative to the amount of content submitted) tend to have more more positive sentiment in their discussion of the moderators, and less negative sentiment. Communities with fewer than five posts and comments per mod per day use 2.5×2.5$\times$2.5 × as much positive sentiment in their mod discourse compared to communities with more than 100 posts and comments per mod per day.
5.1 Content Removal and Moderator Workload
@@
Method. We can identify content removed by moderators in each community by counting the occurrences of ‘[removed]’ posts and comments within each community. By dividing these counts by the total amount of content submitted to that community, we can compute the total percentage of content removed by mods. Using our moderator timelines (§3.1), we can compute the total number of moderator-tenures in any given subreddit over the course of our analysis time period. We also estimate the expected ‘workload’ of each mod by dividing the total daily amount of content submitted to each community by the number of mods available to review that content; while in reality it is unlikely that all moderators share the work of reviewing content evenly (Li, Hecht, and Chancellor 2022a); this metric nonetheless helps us understand the ratio of moderators to content within each community. If workload was shared unevenly, this would only make the workload even higher for a subset of moderators, leading to at worst to overly conservative estimates (higher workload to be addressed by fewer moderators).
@@
Results. We find that communities with higher mod workloads are associated with a decreasing fraction of negative mod discourse (Figure 5). Communities with fewer than five pieces of daily content for each mod (8.8% of all communities) have more mod discourse with positive sentiment than with negative sentiment, a rarity on a platform where mods are far more commonly discussed negatively. As the mod workload increases and the number of mods decreases relative to the amount of content in the community, the fraction of positive mod discourse decreases significantly, with communities with more than 100 posts and comments per mod per day using positive sentiment to describe their mods only 0.39×0.39$\times$0.39 × as often as the 8.8% of communities with the lowest mod workload.
@@
Generally, we find that the fraction of mod discourse with negative sentiment is higher in communities with more removed content (Appendix Figure 10). However, these trends vary depending on the topic of the community (Figure 6). [^5] Similar figures for each of the six community topics are included in Appendix Figure 11. Compared to all other communities, news communities seem to reverse this trend, with perceptions of moderators improving as the amount of content removed from news communities increases. Amongst news communities that remove $<0.5\%$ of their content, 60% of mod discourse has negative sentiment, which drops 35% to 39% for news communities whose mods remove between 1.5% and 3% of their content. For the same amounts of removed content for non-news communities, the fraction of mod discourse with negative sentiment jumps 32%, from 24% to 32%.
@@
Figure 6: For most communities, the fraction of mod discourse with negative sentiment increases as community moderators remove more content (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment actually dropping 35% from news communities whose mods remove $<0.5\%$ of content to communities whose mods remove between 1.5% and 3% of content (b).
Figure 7: Newly appointed mods are associated with a greater improvement in mod perceptions if they are engaged in the community and elsewhere on reddit before their tenure, and if they are engaged during their tenure. (a-b) On average, adding a moderator to a community results in a decrease in the fraction of mod discourse in that community with negative sentiment (red), regardless of the degree to which that moderator has engaged with the community prior to, or will engage with during, their tenure. However, adding a moderator who already has or will engage with the community also results in an increase in the fraction of mod discourse in the community with positive sentiment (aqua), and that increase is 32.5% larger when adding a mod who will engage with the community going forward (b) than for one who already has (a). Adding a moderator who is an active member of communities other than the one they are becoming a mod of results in an increase in positive, and a decrease in negative, sentiment in mod discourse.
@@
Implications. Our results suggest that adding additional moderators to a community (and therefore reducing the effective moderator workload) may improve community members’ perceptions of their mods (Figure 5), although our correlational results may or may not imply causation. In §5.2 we report on additional longitudinal evidence that recruiting additional moderators can have a positive effect. However, simply increasing the amount of content which is removed does not, in general, appear to be associated with more positive moderator discourse sentiment—in fact, the opposite appears to be the case for communities that are not focused on sharing news (Figure 6). Taken together, these results imply that there are topic-specific nuances to content removal, and that mods should use care when deciding to remove more or less content, e.g., by increasing or decreasing the strictness by which rules are enforced. Our results suggest that certain topics are more amenable to stricter rule enforcement than others. While the members of news communities feel less negatively about their mods as the mods remove more content, members of discussion communities whose mods remove $\approx 1.5\%$ of content use 1.75×1.75$\times$1.75 × as much negative-sentiment language to discuss their mods as members of discussion communities whose mods remove $<0.5\%$ of content–a steep increase (Appendix Figure 11).
5.2 Community Engagement
@@
Method. Some moderators are actively engaged with community, soliciting feedback from non-moderators, updating community members on moderation-related news, and contributing to regular community content in addition to their official moderator duties. Other moderators are far less visible, opting to remove content and change rules without participating in the community more broadly. Furthermore, many moderators also participate in other communities beyond just the one(s) they moderate. Using moderators’ public posts and comments, for each moderator-appointment to a community, we compute the number of posts and comments they made in that community and in other communities before and during their tenure as a moderator. We combine these counts with the change in the fraction of mod discourse immediately before and after each mod’s appointment (§3.3) to associate the impact that each mod had on the community to which they were appointed.
@@
Results. We find that, regardless of the moderators’ previous engagement (or lack thereof) with the community, adding a new moderator to a community results in a decrease in the fraction of mod discourse with negative sentiment (Figure 7). However, not all moderators have the same impact. The most positively impactful moderators are those who are engaged with the community prior to their appointment (Figure 7a), will continue to engage with the community during their tenures (Figure 7b), and were also active in other communities prior to their appointment (Figure 7c). New mods who do all three of these things are associated with a 2.5pp increase in the fraction of positive mod discourse, whereas mods who do none of these things are instead associated with a 0.5pp decrease.
@@
Implications. Our results suggest that moderators who actively engage with the community both before and during their tenures are most likely to have a positive impact on the community’s perception of their mod team, which may be because these moderators are more familiar with the community, and are better mods as a result. Another plausible mechanism for this effect is that non-moderator community members value the transparency and accountability that may stem from increased moderator engagement. Our results also suggest that moderators who are active in other subreddits beyond the one(s) they moderate have a more positive impact on the communities they moderate, perhaps because participating in a broader range of communities makes them more effective moderators (Zhu, Kraut, and Kittur 2014). Lastly, our results suggest that of these factors, engagement with the community during the moderator’s tenure has the largest impact.
5.3 Moderator Recruiting
@@
Figure 8: Communities that recruit moderators publicly are 8.78×8.78$\times$8.78 × larger than the average community which recruits only privately, in terms of the community’s daily volume of content (a). Small communities lean towards private recruiting. (b) Compared to private recruiting, recruiting moderators publicly is more polarizing, with an increase in both positive and negative fractions of mod discourse, and a corresponding decrease in neutral sentiment.
@@
Implications. Our findings suggest that public moderator recruiting has the potential be a powerful tool to improve community perceptions of moderators when used carefully, and can be harmful when used without regard to the community’s preferences. A plausible explanation for the increased polarization resulting from public recruiting is that in some circumstances, the public nature of the recruiting exacerbates existing frustrations with mod teams, for example if due process was not followed during moderator elections, or if a perceived-outside was brought in via external recruiting when community members themselves preferred someone with more experience in the community. More work is needed to disentangle confounding factors and more deeply assess the differences between different moderators’ appointments.
5.4 Novice and Experienced Mods

### v2

@@ 4. How are moderators of different communities perceived differently by their communities?
Method. To quantify a community’s health, we divide communities into quartiles based on their community members’ responses to a recent survey (§3.2). Survey responses were collected between May-July 2021, a period overlapping the the end of our mod discourse data time range. For this analysis only, we exclude communities which were not surveyed. Community topic is classified into one of six topical categories using our topic classifier (§3.2). We quantify community size as the volume of submitted content: the average number of posts and comments per day over our study period, which we use to group communities with similar size. For each group, we compute the amount of mod discourse as well as the composition of that discourse.
@@
Results. Different aspects of community health are associated with better and worse perceptions of moderators. The smallest-feeling 25% of communities, based on member self-reports, use 27% more positive and 16% less negative sentiment to discuss their moderators than the largest-feeling 25% of communities (Figure 2a). Figure 2a-e shows that communities which rate themselves as having higher quality content and being more trustworthy, more inclusive, and more safe all use more positive and less negative to discuss their moderators, as well.
Figure 3. Perceptions of moderators vary significantly across communities with different sizes (a-b) and topics (c-d). In general, smaller communities devote a relatively larger proportion of their content to discussing their moderators (a), and smaller communities express more positive and less negative sentiment towards their mods (b). Discussion, meme-sharing, and news communities have proportionally more mod discourse (c), while meme and news-sharing communities exhibit the most negative sentiment towards their moderators (d).
@@
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities have more mod discourse (Figure 3a), with tiny communities with $<10$ posts+comments/day having 7.6×7.6$\times$7.6 × as many posts and comments discussing moderators as huge communities (those with more than 10k posts and comments per day), relative to the total amount of content. Tiny communities also have 6.0×6.0$\times$6.0 × more positive mod discourse, and 0.46×0.46$\times$0.46 × as much negative discourse, than huge communities.
@@
Examining communities of different topics, discussion communities have the most mod discourse, on average (Figure 3c). Hobby communities have generally the most positive mod discourse, with meme and news-sharing communities having 46% more of their mod discourse have negative sentiment than communities of other topics (Figure 3d).
@@
Implications. Surveys of redditors have shown that they consider a wide range of factors to be important to the overall ‘health’ of their communities (Weld et al., 2024). Here we show that communities that are doing well with regards to factors widely considered to be important, such as safety and quality of content, tend to perceive their moderators more positively. Many of these factors are only indirectly controlled by the moderators, for example quality of content—while moderators can set and enforce rules aimed at improve the quality of content in their communities, quality of content is ultimately a function of the content submitted by community members, not moderators. This can lead to moderators being ‘blamed’ for problems largely outside of their control, which has been previously shown in small-n surveys (Matias, 2019a; Jhaver et al., 2023a) and is further supported empirically here, as shown in Figure 2.
@@
Small communities appear to both discuss their moderators more, as well as use more positive sentiment in their discussions (Figure 3a,b). Several factors may contribute to this, including that the increased anonymity that comes with participating in a larger community may make people feel more comfortable speaking negatively about the moderators, and that smaller communities are more likely to be newly formed and thus still establishing moderation norms, leading to more mod discourse (Hwang and Foote, 2021; Seering and Kairam, 2022).
5. Adjusting for Confounders using IPTW & DID
While we are ideally interested in causal, actionable insights, our study is a retrospective observational study, and like any other, subject to potential confounding. In §4 we identified two important confounders, size and topic. Next we describe two methods from the causal inference literature which we use to adjust for these and other confounding factors in subsequent analyses.
Inverse Probability of Treatment Weighting (IPTW). IPTW is a statistical method to adjust for imbalance in possible confounding factors when making comparisons between different observed groups (e.g., a treatment and a control group). We select IPTW over other methods, such as stratification, due to its excellent efficiency (Austin and Stuart, 2015). In §6, we compute the probability of treatment (propensity score, $P(Z|\mathbf{X})$) by applying a logistic regression model to community covariates $\mathbf{X}$ including the topic and size of the community. A complete list of covariates for each analysis is given in Supplementary Materials  C. We generalize IPTW to non-binary treatments to estimate a dose-response curve using the method from (Althoff et al., 2016) in which the treatment is discretized into bins and a different propensity score is computed for each treatment bin in a ‘one-versus-rest’ scheme. For final analyses, each observation is weighted by the inverse of the probability of the treatment it received, such that the weight for observation $i$ is $w_{i}=\frac{Z_{i}}{P(Z_{i}=1|\mathbf{X}_{i})}+\frac{1-Z_{i}}{P(Z_{i}=0|\mathbf{X}_{i})}$, where $Z_{i}$ is the treatment received by observation $i$ (0 for control, 1 for treated), and $\mathbf{X}_{i}$ is a vector of the covariates. In analyses where we use IPTW, we also include the non-IPTW adjusted estimate in gray. In these analyses, reweighting does not dramatically impact our findings, suggesting that the impact of these confounders is moderate at best.
An important validity check for IPTW is to assess the balance of covariates for each treatment group after weighting as applied (Austin and Stuart, 2015). Two groups are often considered ‘balanced’ or ‘indistinguishable‘ if all covariates are within a standardized mean difference (SMD) of 0.25 standard deviations (Althoff et al., 2016). For each treatment group, we compute the difference between each covariate’s weighted mean value and the reference distribution, consisting of the entire population. We compute the standardized mean difference (SMD) by normalizing the difference in means by the standard deviation of the values of the reference distribution. Our reweighting method achieves balance for every covariate substantially related to the treatment. SMDs after weighting are given for each covariate and each analysis in Supplementary Materials C.
Difference in Difference Analyses (DID). Difference in difference analyses enable the estimation of the impact of an intervention by comparing the outcome before vs. after the intervention was applied to the treated group, and comparing this before vs. after difference to an untreated control group. For our analyses of the impact of moderators’ community engagement (§6.2), we use modified DID analyses with multiple time periods (Callaway and Sant’Anna, 2021) to compare the difference in communities’ receptions of new moderators who are engaged with the community (the ‘treatment’) vs. those who are not (the ‘control’). Two four week long periods immediately preceding and following each moderators’ appointment were used to compute the first-order difference. Since mod discourse naturally varies over time, averaging over many DIDs from many new moderator appointments serves to reduce confounding background temporal trends by comparing the difference to an untreated control group.
6. What can moderators do to improve how they are perceived by their community?
@@
Moderators have a great deal of autonomy to run their communities as they see fit, including enforcing rules by removing content and growing the mod team by recruiting or appointing new moderators. Moderators also may (or may not) participate in the community as ‘regular community members’ in addition to their mod duties. In this section, we identify promising suggestions for moderators by comparing communities whose moderators have different workloads, rule enforcement strategies, and degrees of community engagement.
6.1. Content Removal and Moderator Workload
@@
Method. We can identify content removed by moderators in each community by counting the occurrences of ‘[removed]’ posts and comments within each community. By dividing by the total amount of content submitted to that community, we can compute the total percentage of content removed by mods. Using our mod timelines (§3.1), we can compute the total number of moderator-tenures in any given subreddit over the course of our analysis time period. We also approximate ‘workload’ of each mod by dividing the total amount of content submitted to each community by the number of mods available to review that content; while in reality it is unlikely that all moderators share the work of reviewing content evenly (Li et al., 2022a); this metric nonetheless helps us understand the ratio of moderators to content within each community. If workload was shared unevenly, this would only make the workload even higher for a subset of moderators, leading to at worst to overly conservative estimates (higher workload to be addressed by fewer moderators).
@@
Results. We find that communities with higher mod workloads have more negative mod discourse (Figure 4). Communities with fewer than five pieces of daily content for each mod (8.8% of all communities) have approximately equal amounts of mod discourse with positive and negative sentiment, a rarity on a platform where mods are far more commonly discussed negatively. In communities with higher mod workload, the composition of mod discourse is much less positive, with communities with more than 100 posts and comments per mod per day using positive sentiment to describe their mods only 0.39×0.39$\times$0.39 × as often as the 8.8% of communities with the lowest mod workload.
Figure 4. Moderators in communities with lower workloads are perceived more positively and less negatively than moderators in communities with high workloads. Communities with lower moderator workloads (more moderators relative to the amount of content submitted) tend to have more more positive sentiment in their discussion of the moderators, and less negative sentiment. Communities with fewer than five posts and comments per mod per day use 2.5×2.5$\times$2.5 × as much positive sentiment in their mod discourse compared to communities with more than 100 posts and comments per mod per day.
@@
Generally, we find that the composition of mod discourse is more negative in communities with more removed content (Supplementary Figure 9). However, these trends vary depending on the topic of the community, with news communities in particular showing the opposite trend (Figure 5). [^2] Similar figures for each of the six community topics are included in Supplementary Figure 10. Amongst news communities that remove $\leq 1\%$ of their content, 48% of mod discourse has negative sentiment, which drops 11pp to 39% for news communities whose mods remove between 2% and 3% of their content. For the same amounts of removed content for non-news communities, the fraction of mod discourse with negative sentiment increases 6pp, from 26% to 33%.
@@
Figure 5. For most topics, communities where moderators remove more content exhibit more negative sentiment (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment 11pp lower in news communities whose mods remove less than $1\%$ of content compared to communities whose mods remove 2% - 3% of content (b).
@@
Implications. Our results suggest that adding additional moderators to a community (and therefore reducing the effective moderator workload) may improve community members’ perceptions of their mods (Figure 4). In §6.2 we report on additional longitudinal evidence that recruiting additional moderators can have a positive effect. However, simply increasing the amount of content which is removed does not, in general, appear to be associated with more positive moderator discourse sentiment—in fact, the opposite appears to be the case for communities that are not focused on sharing news (Figure 5). Taken together, these results imply that there are topic-specific nuances to content removal, and that mods should use care when deciding how strictly to enforce rules, and how much content to remove. Our results suggest that certain topics are more amenable to stricter rule enforcement than others.
6.2. Community Engagement
@@
Method. Some moderators are actively engaged with community, soliciting feedback from non-moderators, updating community members on moderation-related news, and contributing to regular community content in addition to their official moderator duties. Other moderators are far less visible, opting to remove content and change rules without participating in the community more broadly. Furthermore, many moderators also participate in other communities beyond just the one(s) they moderate. Using moderators’ public posts and comments, for each moderator-appointment to a community, we compute the number of posts and comments they made in that community and in other communities before and during their tenure as a moderator. We use DID analyses (§5) to estimate the impact of appointing new mods with different degrees of community engagement (or lack thereof).
@@
Results. We find that, regardless of the moderators’ previous engagement with the community, adding a new moderator to a community is associated with a decrease in mod discourse with negative sentiment (Figure 6). However, not all moderators have the same impact. The most positively impactful moderators are those who are engaged with the community prior to their appointment (Figure 6a), will continue to engage with the community during their tenures (Figure 6b), and were also active in other communities prior to their appointment (Figure 6c). New mods who do all three of these things are associated with a 2.5pp increase in positive mod discourse, whereas mods who do none of these things are associated with a 0.5pp decrease.
@@
Implications. Our results suggest that moderators who actively engage with the community both before and during their tenures are most likely to have a positive impact on the community’s perception of their mod team, which may be because these moderators are more familiar with the community, and are better mods as a result. Another plausible mechanism for this effect is that non-moderator community members value the transparency and accountability that may stem from increased moderator engagement. Our results also suggest that moderators who are active in other subreddits beyond the one(s) they moderate have a more positive impact on the communities they moderate, perhaps because participating in a broader range of communities makes them more effective moderators (Zhu et al., 2014). Lastly, our results suggest that of these factors, engagement with the community during the moderator’s tenure has the largest impact.
Figure 6. Newly appointed mods are associated with a greater improvement in mod perceptions if they are engaged in the community and elsewhere on reddit before their tenure, and if they are engaged during their tenure (a-c). Adding a moderator who already has or will engage with the community is associated with in an increase in the fraction of mod discourse in the community with positive sentiment (aqua), and that increase is 32.5% larger when adding a mod who will engage with the community going forward (b) than for one who already has (a). Adding a moderator who is an active member of communities other than the one they are becoming a mod of is associated with an increase in positive, and a decrease in negative, sentiment in mod discourse (c).
6.3. Moderator Recruiting
@@
Figure 7. Communities that recruit moderators publicly are 8.78×8.78$\times$8.78 × larger than the average community which recruits only privately, in terms of the community’s daily volume of content (a). Small communities lean towards private recruiting. (b) Compared to private recruiting, recruiting moderators publicly is polarizing: it is associated with an increase in both positive and negative fractions of mod discourse, and a corresponding decrease in neutral sentiment.
@@
Implications. Our findings suggest that public moderator recruiting has the potential be a powerful tool to improve community perceptions of moderators when used carefully, and can be harmful when used without regard to the community’s preferences. A plausible explanation for the increased polarization resulting from public recruiting is that in some circumstances, the public nature of the recruiting exacerbates existing frustrations with mod teams, for example if due process was not followed during moderator elections, or if a perceived-outside was brought in via external recruiting when community members themselves preferred someone with more experience in the community. More work is needed to assess the differences between different moderator recruiting strategies.
6.4. Novice and Experienced Mods